OpenStudy (anonymous):

How many different four person committees can be formed from a group of six boys and four girls?

OpenStudy (anonymous):

the answer is 10

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

the answer is most assuredly not ten

Parth (parthkohli):

But is it mentioned that we'll have 2 boys and 2 girls?

Parth (parthkohli):

It'd be 10c4 according to me.

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow _{10}C_{4} = {10! \over (10 - 4)! \times 4!} }\)

OpenStudy (anonymous):

and my answer is also wrong, because i thought it was two boys and two girls. wrongo answer is \[\dbinom{10}{4}=\frac{10\times 9\times 8\times 7}{4\times 3\times 2}=10\times 3\times 7\]

Parth (parthkohli):

Thank you satellite!